40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5115     Entry time: Thu Aug 4 01:49:08 2011
Author: Suresh 
Type: Update 
Category: IOO 
Subject: MC is locked 

I measured the power transmitted from the PSL to the MC. It is 19mW.

The MC is now locked.  The MC Autolocker script cannot be used now since the tigger conditions are not met.  It has been disabled on the C1IOO-LOCK_MC screen.  The boost switch also is set to zero.  Increasing the boost results in MC unlocking. 

The C1:IOO-MC_RFPC_DCMON was going from 1.4 (MC Unlocked) to 0.66 (MC_locked).  I thought we ought to have a factor of ten drop in this since under high power conditions we used to have a drop of about 5.6 to 0.6.   So I adjusted the zig-zag at the end of the PSL table to improve the alignment.  It now goes from 1.4 to 0.13 when the MC is locked.   The lock is also much more stable now.  It still does not tolerate any boost though.

I checked to make sure that the beam centering on MC_REFL PD is optimal since I touched the zig-zag.  The RFPD output is now 0.7V (MC unlocked).  This matches well with the fact that we used to have 3.5V on it with the MC unlocked.  And we have cut the down the power incident on this by a factor of 5.  Because 1W -> 20mW at the PSL table  and 10% BS -> 100% Y1-...

 

 

ELOG V3.1.3-